# Open dynamical systems with a moving hole

### v1

@@ 1.1. Open dynamical systems with a moving hole
    $\dim_{H}\left\{\omega\in\left(D_{b}^{m}\right)^{\mathbb{N}}:\dim_{H}K^{\omega}=\alpha,\leavevmode\nobreak\ \dim_{P}K^{\omega}=\beta\right\}$ $=\min\left\{\dim_{H}L_{\alpha},\dim_{H}L_{\beta}\right\},$
    $\dim_{P}\left\{\omega\in\left(D_{b}^{m}\right)^{\mathbb{N}}:\dim_{H}K^{\omega}=\alpha,\leavevmode\nobreak\ \dim_{P}K^{\omega}=\beta\right\}$ $=\max\left\{\dim_{H}L_{\alpha},\dim_{H}L_{\beta}\right\}.$
@@
    In the proof of Theorem 1.6 we obtain the following estimation on the size of $\|A_{\omega^{0}}\cdots A_{\omega^{n-1}}\|$. For any $\omega\in\left(D_{b}^{m}\right)^{\mathbb{N}}$ there exists a constant $C\geq 1$ such that for all $n\in\mathbb{N}$,
@@ 1.2. Badly approximable numbers and non-recurrence points
Given an integer $b\geq 3$, let $T_{b}:[0,1)\to[0,1);\leavevmode\nobreak\ x\mapsto bx\pmod{1}$ be the expanding map on the circle. For a sequence of open balls $\{B_{n}\}_{n=0}^{\infty}$ in $[0,1)$, the following set
@@
$\begin{split}K\left(\left\{B_{n}\right\}\right)&:=[0,1)\setminus W(\left\{B_{n}\right\})=\left\{x\in[0,1):T_{b}^{n}x\notin B_{n}\textrm{ for all but finitely many }n\right\}\end{split}$
@@
Suppose the limit $\lim_{n\to\infty}\operatorname{diam}(B_{n})$ exists. Then
$\dim_{H}K\left(\left\{B_{n}\right\}\right)=1\quad\Longleftrightarrow\quad\lim_{n\to\infty}\operatorname{diam}(B_{n})=0.$
@@
    The assumption on the existence of $\lim_{n\to\infty}\operatorname{diam}(B_{n})$ can not be removed. Even if $\liminf_{n\to\infty}\operatorname{diam}(B_{n})=0$, we can still have $\dim_{H}K\left(\left\{B_{n}\right\}\right)<1$. For example, let $\operatorname{diam}(B_{2n-2})=n^{-1}$ and $B_{2n-1}=[0,1)$ for all $n\in\mathbb{N}$. Then $\liminf_{n\to\infty}\operatorname{diam}(B_{n})=0$. So, for all $x\in[0,1)$ we have $T^{2n-1}_{b}x\in[0,1)=B_{2n-1}$ for all $n\in\mathbb{N}$. This implies $K\left(\left\{B_{n}\right\}\right)=\emptyset$, and thus $\dim_{H}K\left(\left\{B_{n}\right\}\right)=0$. On the other hand, if $\dim_{H}K\left(\left\{B_{n}\right\}\right)=1$ then we can deduce $\liminf_{n\to\infty}\operatorname{diam}(B_{n})=0$.
@@
$R(\phi):=\left\{x\in[0,1):|T_{b}^{n}x-x|<\phi(n)\textrm{ for infinitely many }n\right\}$
@@
$E\left(\phi\right):=[0,1)\setminus R(\phi)=\left\{x\in[0,1):|T_{b}^{n}x-x|\geq\phi(n)\textrm{ for all but finitely many }n\right\}.$
@@
Suppose the limit $\lim_{n\to\infty}\phi(n)$ exists. Then
$\dim_{H}E\left(\phi\right)=1\quad\Longleftrightarrow\quad\lim_{n\to\infty}\phi(n)=0.$
@@
Similar to Remark 1.9, the assumption on the existence of the limit $\lim_{n\to\infty}\phi(n)$ can not be removed.
1.3. Finiteness property of joint spectral radius
Given a finite set of $d\times d$ real matrices =$\mathcal{M}=\left\{M_{1},\ldots,M_{r}\right\}$, Rota and Strang [44] introduced the concept of joint spectral radius
(1.6) $\hat{\rho}(\mathcal{M}):=\lim_{n\to\infty}\sup_{i_{1}\ldots i_{n}\in\left\{1,2,\ldots,r\right\}^{n}}\left\|M_{i_{1}}\cdots M_{i_{n}}\right\|^{1/n},$
@@ 1.3. Finiteness property of joint spectral radius
$\hat{\rho}(\mathcal{M})=\rho\left(M_{i_{1}}\cdots M_{i_{n}}\right)^{1/n},$
where $\rho(M)$ denotes the spectral radius of a matrix $M$. Lagarias and Wang [35] conjectured that every finite set of $d\times d$ real matrices possesses the finiteness property. However, this conjecture was proven to be false in a general setting (cf. [6, 8, 28]); but it still holds in many situations (cf. [1, 13, 30, 34, 39, 42]). In particular, if $\mathcal{M}$ is a finite set of $d\times d$ nonnegative matrices with $\hat{\rho}(\mathcal{M})=\rho(\sum_{M\in\mathcal{M}}M)$, then $\mathcal{M}$ has the finiteness property (cf. [14]).
Jungers and Blondel [31, Theorem 4] established the following key equivalence: the finiteness property holds for all sets of nonnegative rational square matrices if and only if it holds for all pairs of binary square matrices. They also proved that the finiteness property holds for all pairs of $2\times 2$ binary matrices. Note that each adjacency matrix $A_{\mathbf{d}}$ of the directed graph $G_{\mathbf{d}}$ is a $0-1$ matrix. Our final result says that the set $\left\{A_{\mathbf{d}}:\mathbf{d}\in D_{b}^{m}\right\}$ of adjacency matrices satisfies the finiteness property.
@@
The set $\mathcal{A}=\left\{A_{\mathbf{d}}:\mathbf{d}\in D_{b}^{m}\right\}$ has the finiteness property. Furthermore, for any periodic and progressively overlapping sequence $\omega=(\omega^{0}\omega^{1}\ldots\omega^{n-1})^{\infty}\in\left(D_{b}^{m}\right)^{\mathbb{N}}$ we have
$\hat{\rho}(\mathcal{A})=\rho(A_{\omega^{0}}A_{\omega^{1}}\cdots A_{\omega^{n-1}})^{1/n}=\lambda,$
@@
Indeed Theorem 1.12 shows that the finiteness property $\hat{\rho}(\mathcal{A})=\rho(A_{\omega^{0}}A_{\omega^{1}}\ldots A_{\omega^{n-1}})^{1/n}$ holds for infinitely many blocks $\omega^{0}\omega^{1}\ldots\omega^{n-1}\in\left(D_{b}^{m}\right)^{*},n\in\mathbb{N}$.
The rest of the paper is organized as follows. In the next section we consider the dimensions of $K^{\omega}$ for a general $\omega\in(D_{b}^{m})^{\mathbb{N}}$, and prove Theorem 1.1. When $\omega$ is progressively overlapping or totally distinct, we show in Section 3 that all of these fractal dimensions of $K^{\omega}$ are equal. In fact, we show in this case that $K^{\omega}$ is Ahlfors regular (see Theorem 1.4). In Section 4, we prove the intermediate value property for the dimensions of $K^{\omega}$, and show that there are infinitely many $\omega\in(D_{b}^{m})^{\mathbb{N}}$ such that the dimensions of $K^{\omega}$ are not equal (see Theorem 1.6). As applications, we prove in Section 5 Theorems 1.8 and 1.10 for badly approximable numbers in Diophantine approximation and non-recurrence points, respectively; and prove the finiteness property for joint spectral radius of matrices (see Theorem 1.12). Finally, in Section 6 we make some remarks on possible extensions of our work to higher dimensions or to homogeneous self-similar IFSs.
@@ 2. Hausdorff and packing dimensions of $K^{\omega}$
In this section we will determine the Hausdorff and packing dimensions of $K^{\omega}$ for a general sequence $\omega\in(D_{b}^{m})^{\mathbb{N}}$, and prove Theorem 1.1. Note that $b\in\mathbb{N}_{\geq 3}$ and $m\in\mathbb{N}_{\geq 2}$. For a sequence $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$, recall the definitions of $\Sigma^{\omega}$ and $\Sigma_{k}^{\omega}$ from (1.1) and (1.3), and one can easily check that
(2.1) $\Sigma_{k}^{\omega}=\begin{cases}D_{b}^{k},&\mbox{if }1\leq k<m,\\ \big{\{}d_{1}d_{2}\ldots d_{k}\in D_{b}^{k}:d_{i+1}\ldots d_{i+m}\neq\omega^{i}\leavevmode\nobreak\ \forall 0\leq i\leq k-m\big{\}},&\mbox{if }k\geq m.\end{cases}$
@@
For any $\omega\in(D_{b}^{m})^{\mathbb{N}}$ and for any $k\in\mathbb{N}$ we have
@@
Note by (2.1) that $|\Sigma_{k}^{\omega}|=b^{k}$ for all $1\leq k<m$ and $|\Sigma_{m}^{\omega}|=b^{m}-1$. Thus the inequalities hold for all $1\leq k<m$. Now we assume $k\geq m$, and write $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ with each $\omega^{i}=\omega^{i}_{1}\omega^{i}_{2}\ldots\omega^{i}_{m}\in D_{b}^{m}$. If $d_{1}d_{2}\ldots d_{k}\in\Sigma_{k}^{\omega}$ and $d_{k+1}\in D_{b}\setminus\left\{\omega_{m}^{k-m+1}\right\}$, then by (2.1) we must have $d_{1}d_{2}\ldots d_{k}d_{k+1}\in\Sigma_{k+1}^{\omega}$. So, we obtain $|\Sigma_{k+1}^{\omega}|\geq(b-1)|\Sigma_{k}^{\omega}|$. On the other hand, if $d_{1}d_{2}\ldots d_{k+1}\in\Sigma_{k+1}^{\omega}$, then by (2.1) its prefix $d_{1}d_{2}\ldots d_{k}\in\Sigma_{k}^{\omega}$. Clearly, we have $d_{k+1}\in D_{b}$. It follows that $|\Sigma_{k+1}^{\omega}|\leq b|\Sigma_{k}^{\omega}|$. ∎
@@
Let $\omega\in(D_{b}^{m})^{\mathbb{N}}$ and $k\in\mathbb{N}_{\geq m}$. Then for $\mathbf{c}\in\Sigma_{k}^{\omega}$ and $n\in\mathbb{N}$ we have
@@
Write $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ with each $\omega^{i}=\omega^{i}_{1}\omega^{i}_{2}\ldots\omega^{i}_{m}\in D_{b}^{m}$. Since $b\geq 3$, we can choose a word $u_{1}\ldots u_{m-1}\in D_{b}^{m-1}$ such that
@@
Let $\tau:=\omega^{k+m-1}\omega^{k+m}\ldots\in(D_{b}^{m})^{\mathbb{N}}$. We claim that
$\mathbf{c}u_{1}\ldots u_{m-1}\mathbf{d}\in\Sigma_{k+n}^{\omega}\quad\text{for any}\leavevmode\nobreak\ \mathbf{d}\in\Sigma_{n-m+1}^{\tau}.$
@@
$\Sigma_{k+n}^{\omega}=\bigcup_{\widetilde{\mathbf{c}}\in\Sigma_{k}^{\omega}}\Sigma_{k+n}^{\omega}(\widetilde{\mathbf{c}}),$
@@
For $\mathbf{d}\in D_{b}^{m}$, recall that $A_{\mathbf{d}}$ is the adjacency matrix of the directed graph $G_{\mathbf{d}}$, which describes the shift of finite type $X_{\mathbf{d}}$. Note that for a matrix $A=(a_{i,j})$ we use the norm $\|A\|=\sum_{i,j}|a_{i,j}|$.
@@
For any $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ and $k\in\mathbb{N}_{\geq m}$ we have
@@
$\prod_{i=0}^{k-m}a_{i}\big{(}d_{i+1}\ldots d_{i+m-1},\leavevmode\nobreak\ d_{i+2}\ldots d_{i+m}\big{)}.$
@@
Since $k_{i}\geq k_{0}$, by (2.8) and (2.10) we have $|\Sigma_{k_{i}}^{\omega}|^{-1}\leq b^{-k_{i}t}<b^{s}\operatorname{diam}(U_{i})^{t}$. Then (2.9) follows directly from (2.12). Thus we obtain $\dim_{H}K^{\omega}\geq t$. Since $t\in(0,s)$ was chosen arbitrarily, we conclude that $\dim_{H}K^{\omega}\geq s$.

### v2

@@ 1.1. Open dynamical systems with a moving hole
    $\dim_{H}\left\{\omega\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}}:\dim_{H}K^{\omega}=\alpha,\penalty 10000\ \dim_{P}K^{\omega}=\beta\right\}$ $=\min\left\{\dim_{H}L_{\alpha},\dim_{H}L_{\beta}\right\},$
    $\dim_{P}\left\{\omega\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}}:\dim_{H}K^{\omega}=\alpha,\penalty 10000\ \dim_{P}K^{\omega}=\beta\right\}$ $=\max\left\{\dim_{H}L_{\alpha},\dim_{H}L_{\beta}\right\}.$
@@
    In the proof of Theorem 1.6 we obtain the following estimation on the size of $\|A_{\omega^{0}}\cdots A_{\omega^{n-1}}\|$. For any $\omega\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}}$ there exists a constant $C\geq 1$ such that for all $n\in\mathbb{N}$,
@@
Finally, we show that if the progressively overlapping pattern and the totally distinct pattern occur periodically in the sequence $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$, then the set $K^{\omega}$ is Ahlfors regular. More precisely, for $p,q\in\mathbb{N}$ let $L_{p,q}$ be the set of sequences $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ that is progressively overlapping at position-$k$ for $\ell(p+q)<k\leq\ell(p+q)+p$ and is totally distinct at position-$k$ for $\ell(p+q)+p<k\leq(\ell+1)(p+q)$, where $\ell\in\mathbb{N}_{0}$. In other words, a sequence $\omega=\omega^{0}\omega^{1}\omega^{2}\ldots\in L_{p,q}$ with each $\omega^{k}=\omega^{k}_{1}\omega^{k}_{2}\ldots\omega^{k}_{m}\in D_{b}^{m}$ if it satisfies
$\begin{cases}\omega^{k}_{1}\ldots\omega^{k}_{m-1}=\omega^{k-1}_{2}\ldots\omega^{k-1}_{m},&\mbox{for }\ell(p+q)<k\leq\ell(p+q)+p,\\ \omega^{k}_{1}\neq\omega^{k-1}_{2},\;\omega^{k}_{2}\ldots\omega^{k}_{m-1}=\omega^{k-1}_{3}\ldots\omega^{k-1}_{m}&\mbox{for }\ell(p+q)+p<k\leq(\ell+1)(p+q)\end{cases}$
for all $\ell\in\mathbb{N}_{0}$. Note that for any $(d_{i})\in D_{b}^{\mathbb{N}}$ we can construct a sequence $\omega=\omega^{0}\omega^{1}\omega^{2}\ldots\in L_{p,q}$ by setting $\omega^{0}_{1}\ldots\omega^{0}_{m}=d_{1}\ldots d_{m}$, $\omega^{k}_{1}\ldots\omega^{k}_{m}=d_{k+1}\ldots d_{k+m}$ for $\ell(p+q)<k\leq\ell(p+q)+p$, and $\omega^{k}_{1}\neq d_{k+1},\omega^{k}_{2}\ldots\omega^{k}_{m}=d_{k+2}\ldots d_{k+m}$ for $\ell(p+q)+p<k\leq(\ell+1)(p+q)$, where $\ell\in\mathbb{N}_{0}$. This implies that for any $p,q\in\mathbb{N}$ the set $L_{p,q}$ is uncountable.
Theorem 1.8.
For any $p,q\in\mathbb{N}$, there exists $\lambda_{p,q}>1$ such that the set $K^{\omega}$ is $\frac{\log\lambda_{p,q}}{(p+q)\log b}$-Ahlfors regular for all $\omega\in L_{p,q}$. Moreover, the set
$\bigg\{\frac{\log\lambda_{p,q}}{(p+q)\log b}:p,q\in\mathbb{N}\bigg\}$
is dense in $\frac{\log\eta}{\log b},\frac{\log\lambda}{\log b}]$, where $\lambda$ and $\eta$ are defined as in Theorem 1.4.
@@ 1.2. Badly approximable numbers and non-recurrence points
Given an integer $b\geq 3$, let $T_{b}:[0,1)\to[0,1);\penalty 10000\ x\mapsto bx\pmod{1}$ be the expanding map on the circle. For a sequence of open balls $\{B_{n}\}_{n=0}^{\infty}$ in $[0,1)$, the following set
@@
$\begin{split}K\left(\left\{B_{n}\right\}\right)&:=[0,1)\setminus W(\left\{B_{n}\right\})=\left\{x\in[0,1):T_{b}^{n}(x)\notin B_{n}\textrm{ for all but finitely many }n\right\}\end{split}$
@@
  1. (i)
    If $\lim_{n\to\infty}\operatorname{diam}(B_{n})=0$, then $\dim_{H}K(\left\{B_{n}\right\})=1$.
  2. (ii)
    If $\dim_{H}K(\left\{B_{n}\right\})=1$, then $\liminf_{n\to\infty}\operatorname{diam}(B_{n})=0$.
@@
    If the limit $\lim_{n\to\infty}\operatorname{diam}(B_{n})$ exists, then by Theorem 1.9 it follows that $\dim_{H}K\left(\left\{B_{n}\right\}\right)=1$ if and only if $\lim_{n\to\infty}\operatorname{diam}(B_{n})=0.$
@@
$R(\phi):=\left\{x\in[0,1):|T_{b}^{n}(x)-x|<\phi(n)\textrm{ for infinitely many }n\right\}$
@@
$E\left(\phi\right):=[0,1)\setminus R(\phi)=\left\{x\in[0,1):|T_{b}^{n}(x)-x|\geq\phi(n)\textrm{ for all but finitely many }n\right\}.$
@@
  1. (i)
    If $\lim_{n\to\infty}\phi(n)=0$, then $\dim_{H}E(\phi)=1$.
  2. (ii)
    If $\dim_{H}E(\phi)=1$, then $\liminf_{n\to\infty}\phi(n)=0$.
@@
Similar to Remark 1.10 (i), if the limit $\lim_{n\to\infty}\phi(n)$ exists, then Theorem 1.11 implies that $\dim_{H}E\left(\phi\right)=1$ if and only if $\lim_{n\to\infty}\phi(n)=0.$
1.3. Finiteness property for the joint spectral radius
Given a finite set of $d\times d$ real matrices $\mathcal{M}=\left\{M_{1},\ldots,M_{r}\right\}$, Rota and Strang [44] introduced the concept of the joint spectral radius
(1.6) $\widehat{\rho}(\mathcal{M}):=\lim_{n\to\infty}\sup_{i_{1}\ldots i_{n}\in\left\{1,2,\ldots,r\right\}^{n}}\left\|M_{i_{1}}\cdots M_{i_{n}}\right\|^{1/n},$
@@ 1.3. Finiteness property for the joint spectral radius
$\widehat{\rho}(\mathcal{M})=\rho\left(M_{i_{1}}\cdots M_{i_{n}}\right)^{1/n},$
where $\rho(M)$ denotes the spectral radius of a matrix $M$. Lagarias and Wang [35] conjectured that every finite set of $d\times d$ real matrices possesses the finiteness property. However, this conjecture was proven to be false in a general setting (cf. [6, 8, 28]); but it still holds in many situations (cf. [1, 13, 30, 34, 39, 42]). In particular, if $\mathcal{M}$ is a finite set of $d\times d$ nonnegative matrices with $\widehat{\rho}(\mathcal{M})=\rho(\sum_{M\in\mathcal{M}}M)$, then $\mathcal{M}$ has the finiteness property (cf. [14]).
Jungers and Blondel [31, Theorem 4] established the following key equivalence: the finiteness property holds for all sets of nonnegative rational square matrices if and only if it holds for all pairs of binary square matrices. They also proved that the finiteness property holds for all pairs of $2\times 2$ binary matrices. Note that each adjacency matrix $A_{\mathbf{d}}$ of the directed graph $G_{\mathbf{d}}$ is a 010-1 matrix. Our final result says that the set $\left\{A_{\mathbf{d}}:\mathbf{d}\in D_{b}^{m}\right\}$ of adjacency matrices satisfies the finiteness property.
@@
The set $\mathcal{A}=\left\{A_{\mathbf{d}}:\mathbf{d}\in D_{b}^{m}\right\}$ has the finiteness property. Furthermore, for any periodic and progressively overlapping sequence $\omega=(\omega^{0}\omega^{1}\ldots\omega^{n-1})^{\infty}\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}}$ we have
$\widehat{\rho}(\mathcal{A})=\rho(A_{\omega^{0}}A_{\omega^{1}}\cdots A_{\omega^{n-1}})^{1/n}=\lambda,$
@@
Indeed Theorem 1.13 shows that the finiteness property $\widehat{\rho}(\mathcal{A})=\rho(A_{\omega^{0}}A_{\omega^{1}}\ldots A_{\omega^{n-1}})^{1/n}$ holds for infinitely many blocks $\omega^{0}\omega^{1}\ldots\omega^{n-1}\in\left(D_{b}^{m}\right)^{*},n\in\mathbb{N}$.
The rest of the paper is organized as follows. In the next section we consider the dimensions of $K^{\omega}$ for a general $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$, and prove Theorem 1.1. When $\omega$ is progressively overlapping or totally distinct, we show in Section 3 that all of these fractal dimensions of $K^{\omega}$ are equal. In fact, we show in this case that $K^{\omega}$ is Ahlfors regular (see Theorem 1.4). In Section 4, we prove the intermediate value property for the dimensions of $K^{\omega}$, and show that there are infinitely many $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ such that the dimensions of $K^{\omega}$ are not equal (see Theorem 1.6). In Section 5, we study some special sequences $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ for which $\dim_{H}K^{\omega}=\dim_{P}K^{\omega}$, and prove Theorem 1.8. As applications, we prove in Section 6 Theorems 1.9 and 1.11 for badly approximable numbers in Diophantine approximation and non-recurrence points, respectively; and prove the finiteness property for the joint spectral radius of matrices (see Theorem 1.13). Finally, in Section 7 we make some remarks on possible extensions of our work to higher dimensions or to homogeneous self-similar IFSs.
@@ 2. Hausdorff and packing dimensions of $K^{\omega}$
In this section we will determine the Hausdorff and packing dimensions of $K^{\omega}$ for a general sequence $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$, and prove Theorem 1.1. Note that $b\in\mathbb{N}_{\geq 3}$ and $m\in\mathbb{N}_{\geq 2}$. For a sequence $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$, recall the definitions of $\Sigma^{\omega}$ and $\Sigma_{k}^{\omega}$ from (1.1) and (1.3), and one can easily check that
(2.1) $\Sigma_{k}^{\omega}=\begin{cases}D_{b}^{k},&\mbox{if }1\leq k<m,\\ \big\{d_{1}d_{2}\ldots d_{k}\in D_{b}^{k}:d_{i+1}\ldots d_{i+m}\neq\omega^{i}\penalty 10000\ \forall 0\leq i\leq k-m\big\},&\mbox{if }k\geq m.\end{cases}$
@@
For any $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ and for any $k\in\mathbb{N}$ we have
@@
Note by (2.1) that $|\Sigma_{k}^{\omega}|=b^{k}$ for all $1\leq k<m$ and $|\Sigma_{m}^{\omega}|=b^{m}-1$. Thus the inequalities hold for all $1\leq k<m$. Now we assume $k\geq m$, and write $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ with each $\omega^{i}=\omega^{i}_{1}\omega^{i}_{2}\ldots\omega^{i}_{m}\in D_{b}^{m}$. If $d_{1}d_{2}\ldots d_{k}\in\Sigma_{k}^{\omega}$ and $d_{k+1}\in D_{b}\setminus\left\{\omega_{m}^{k-m+1}\right\}$, then by (2.1) we must have $d_{1}d_{2}\ldots d_{k}d_{k+1}\in\Sigma_{k+1}^{\omega}$. So, we obtain $|\Sigma_{k+1}^{\omega}|\geq(b-1)|\Sigma_{k}^{\omega}|$. On the other hand, if $d_{1}d_{2}\ldots d_{k+1}\in\Sigma_{k+1}^{\omega}$, then by (2.1) its prefix $d_{1}d_{2}\ldots d_{k}\in\Sigma_{k}^{\omega}$. Clearly, we have $d_{k+1}\in D_{b}$. It follows that $|\Sigma_{k+1}^{\omega}|\leq b|\Sigma_{k}^{\omega}|$. ∎
@@
Let $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ and $k\in\mathbb{N}_{\geq m}$. Then for $\mathbf{c}\in\Sigma_{k}^{\omega}$ and $n\in\mathbb{N}$ we have
@@
Write $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ with each $\omega^{i}=\omega^{i}_{1}\omega^{i}_{2}\ldots\omega^{i}_{m}\in D_{b}^{m}$. Since $b\geq 3$, we can choose a word $u_{1}\ldots u_{m-1}\in D_{b}^{m-1}$ such that
@@
Let $\tau:=\omega^{k+m-1}\omega^{k+m}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$. We claim that
$\mathbf{c}u_{1}\ldots u_{m-1}\mathbf{d}\in\Sigma_{k+n}^{\omega}\quad\text{for any}\penalty 10000\ \mathbf{d}\in\Sigma_{n-m+1}^{\tau}.$
@@
[…]
@@
For $\mathbf{d}\in D_{b}^{m}$, recall that $A_{\mathbf{d}}$ is the adjacency matrix of the directed graph $G_{\mathbf{d}}$, which describes the shift of finite type $X_{\mathbf{d}}$. Recall that for a matrix $A=(a_{i,j})$ we use the norm $\|A\|=\sum_{i,j}|a_{i,j}|$.
@@
For any $\omega=\omega^{0}\omega^{1}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ and $k\in\mathbb{N}_{\geq m}$ we have
@@
$\prod_{i=0}^{k-m}a_{i}\big(d_{i+1}\ldots d_{i+m-1},\penalty 10000\ d_{i+2}\ldots d_{i+m}\big).$
@@
Since $k_{i}\geq k_{0}$, by (2.8) and (2.10) we have $|\Sigma_{k_{i}}^{\omega}|^{-1}\leq b^{-k_{i}t}<b^{s}\big(\operatorname{diam}(U_{i})\big)^{t}$. Then (2.9) follows directly from (2.12). Thus we obtain $\dim_{H}K^{\omega}\geq t$. Since $t\in(0,s)$ was chosen arbitrarily, we conclude that $\dim_{H}K^{\omega}\geq s$.